OpenStudy (anonymous):

What are the factors of x2 + 7x + 12?

OpenStudy (goldphenoix):

(x+4)(x+3) You can do the FOIL method to check if it's correct. How did I get this? It's hard to explain. Someone can explain it to you.

OpenStudy (anonymous):

\[x^{2}+7x+12=x^{2}+4x+3x+12=x(x+4)+3(x+4)=(x+4)(x+3)\] Thus (x+4)(x+3) are the required factors of the given expression.

OpenStudy (goldphenoix):

You can only factor a trinomial with two binomial.

OpenStudy (nurali):

x2+7x+12=0 (x+4)(x+3)=0 x=-4,-3

OpenStudy (anonymous):

is it -4,-4 or +4,+3

OpenStudy (goldphenoix):

+4, +3.

OpenStudy (goldphenoix):

Do the FOIL method. (x+4)(x+3) x * x = x^2 x * 3 = 3x 4 * x = 4x 4 * 3 = 12 x^2 + 3x + 4x + 12 Combine like term: 3x + 4x = 7x Therefore, you should get: x^2 + 7x + 12